OpenStudy (anonymous):

Given the system of equations, what is the solution? 3x - 2y + 10 = 0 5y = 4x + 8 {(-34/7, 16/7)} {(34/7, -16/7)} {(-34/7, -16/7)}

OpenStudy (anonymous):

3x - 2y + 10 = 0 5y = 4x + 8 3x - 2y = -10 -4x + 5y = 8 15x - 10y = -50 -8x + 10y = 16 Adding equations: 7x = -34 x = -34/7 = -4.86 5y = 4x + 8 = -11.44 y = -2.29